# Sparse Imagination for Efficient Visual World Model Planning

### v1

@@ Abstract
World model based planning has significantly improved decision-making in complex environments by enabling agents to simulate future states and make informed choices. However, ensuring the prediction accuracy of world models often demands substantial computational resources, posing a major challenge for real-time applications. This computational burden is particularly restrictive in robotics, where resources are severely constrained. To address this limitation, we propose a Sparse Imagination for Efficient Visual World Model Planning, which enhances computational efficiency by reducing the number of tokens processed during forward prediction. Our method leverages a sparsely trained vision-based world model based on transformers with randomized grouped attention strategy, allowing the model to adaptively adjust the number of tokens processed based on the computational resource. By enabling sparse imagination (rollout), our approach significantly accelerates planning while maintaining high control fidelity. Experimental results demonstrate that sparse imagination preserves task performance while dramatically improving inference efficiency, paving the way for the deployment of world models in real-time decision-making scenarios.
@@ 1 Introduction
By “imagining” future trajectories in a learned world model of the environment, agents can perform sophisticated decision making without trial-and-error in the real world hafner2019planet ; hafner2019dreamer . Recent world model advancements for visual control tasks, which necessitate reasoning over high-dimensional image observations, increasingly leverage high-level visual features derived from self-supervised vision models caron2021dino ; oquab2023dinov2 ; nair2022r3m ; xiao2022mvp . For example, Zhou et al. zhou2024dinowm introduces a world model that predicts future Vision Transformer (ViT) patch tokens (DINO features) instead of pixels or single vector representation, enabling zero-shot planning for diverse tasks. By retaining spatially rich visual tokens, such approaches achieve strong generalization on control tasks with complex observations. A downside, however, is that using dozens or hundreds of visual tokens per image incurs high, quadratic computational cost during inference, especially if many rollout simulations are needed for planning each action. This raises a crucial question:
@@
In this paper, we address the above question by introducing the concept of sparse imagination, accelerating world model inference by deliberately using only a sparse subset of visual tokens during forward prediction. Our key insight is that ViT-based image representations contain redundancies raghu2021vision ; pan2021ia ; chen2022principle ; kim2024learning –not all patch tokens are necessary for forward prediction or decision making.
@@
  • We introduce a simple yet effective training and inference method for world models that leverages patch feature dropout to enable sparse visual imagination.
  • Our sparse imagination approach, which utilizes randomly sampling token subsets, achieves substantial inference speedup while preserving competitive task performance across diverse visual control tasks compared to full-patch and single-vector baselines.
  • Our work offers a thorough experimental analysis explaining the operational principles behind the success of the sparse imagination with planning framework.
  • We show through a comparative study that simple random dropout performs comparably to or better than sophisticated token selection and merging methods.
@@ 2 Related Work
World model based decision making, enabling agents to simulate futures for informed choices, has been a long-standing goal in reinforcement learning and robotics kaelbling1996planning ; wm . The field transitioned from computationally heavy pixel-level models finn2016video ; ebert2018visualforesight to latent dynamics models like PlaNet hafner2019planet and Dreamer hafner2019dreamer ; hafner2020dreamerv2 ; hafner2023dreamerv3 , which achieved impressive results by leveraging planning and latent imagination. However, most existing latent world models compress entire images into relatively low-dimensional vectors. While compact, these representations can struggle to capture the fine-grained spatial information necessary for high-precision tasks, as evidenced by their performance on manipulation benchmarks zhou2024dinowm ; tsagkas2025PVR .
@@
A recent trend in decision making leverages powerful pre-trained visual encoders for rich state representations nair2022r3m ; xiao2022mvp . While effective for representation learning, methods using a single image embedding (e.g., CLS token) from encoders like ResNet or ViT often lose crucial fine-grained spatial details, hindering performance on tasks requiring precise spatial understanding, such as manipulation tsagkas2025PVR ; zhou2024dinowm . To address this, approaches using spatial features like ViT patch embeddings for decision making have been developed, exemplified by DINO-WM zhou2024dinowm which uses a set of DINO patch tokens as the state. Building on this paradigm, our work utilizes pre-trained ViT patch features for their spatial granularity. Our key innovation is an inference-time mechanism to drop these tokens during imagination, significantly enhancing computational efficiency. This motivates our central question: how can we retain the effectiveness of rich patch-based representations while substantially reducing computational cost on decision masking?
Figure 1: Sparse Imagination. We propose an accelerated strategy for visual world model based planning with sparse imagination, leveraging a patch token dropout mechanism to achieve efficient model predictive control (MPC) rollouts during inference. At each iteration of MPC, a random subset of visual patches is selected, and predictions are generated exclusively based on these sampled patches. Actions are optimized using the Cross-Entropy Method (CEM) by evaluating their predicted outcomes against the goal in latent space, calculated solely from the selected patches. This random dropout pattern is dynamically updated at every MPC iteration, thereby enhancing the model’s robustness.
@@
The quadratic scaling of ViT self-attention poses a significant computational challenge, particularly for sequential models like world models dosovitskiy2020vit . Improving inference efficiency via token reduction is thus a key vision research area. Existing methods employ diverse strategies such as learned/attention-based selection rao2021dynamicvit ; luo2024ltrp ; yin2022avit ; Liang2021evit ; wang2022vtclfc ; Zhang2023star , merging bolya2022tome ; feng2023efficient ; haurum2024agglomerative , and training-time dropout liu2023patchdropout . These demonstrate efficiency gains from ViT representation redundancy raghu2021vision ; pan2021ia ; chen2022principle ; kim2024learning . In contrast, we present a patch-dropping method specifically for inference-time world model planning. Our approach uniquely applies dropout only during the imagination phase, enabling seamless integration with any transformer world model.
@@ 3 Methodology
We begin by describing the architecture and training procedure of our world model (Section 3.1), followed by introducing sparse imagination, which incorporates patch token dropout for efficient model predictive control (MPC) rollouts (Section 3.2).
@@ 3.2 Sparse Imagination for MPC
For planning with the world model, we utilize MPC using the Cross-Entropy Method (CEM) for trajectory optimization. At each MPC step, candidate action sequences $\{a_{t:t+H-1}\}$ are sampled from a distribution. The trained world model $f_{\theta}$ then predicts future states by iteratively rolling out these action sequences from the current state. In a goal-conditioned setting, candidate sequences are evaluated through an objective such as the MSE loss between the predicted state features and a specified goal state:
$\mathcal{L}_{\text{mpc}}=||\hat{z}_{t+H}-z_{g}||^{2}.$ (2)
@@
The distribution of candidate actions is subsequently updated based on the top-performing sequences. This iterative optimization continues for a predefined iteration step $M$, after which the optimized action sequence is executed in the environment.
@@
However, MPC can be computationally intensive due to significant amount of world model predictions. Specifically, world model rollout occurs $K\times M\times H$ times every MPC step, where the number of candidate sequences is $K$ and planning horizon is $H$. Moreover, this computational cost scales exponentially with the number of visual tokens, making real-time execution increasingly challenging. Since attention layers in the world model are the primary source of this computational burden, reducing the number of tokens processed is crucial for improving MPC efficiency.
@@
We address this problem through sparse imagination. Given that our model is explicitly trained to operate on arbitrary token subsets, we implement random token dropout during MPC planning as shown in Fig. 1. At each MPC step, we randomly generate a token dropout mask characterized by a dropout fraction $p\in[0,1)$ of the user’s choice, denoting the proportion of tokens to discard, keeping randomly sampled $(1-p)N$ tokens. Sparse rollout imagination can effectively reduce redundant computations and improve planning efficiency. Planning a single step might fall to wrong direction when dropping key features but has minimal impact on task performance since it can recover from failures by re-sampling at each planning iteration. The dropout fraction $p$ serves as a controllable parameter balancing computational speed and task accuracy. We empirically evaluate and analyze this trade-off in Section 4.
Figure 2: Randomized Grouped Attention Strategy. , We introduce a randomized grouping strategy during training in order to generalize to arbitrary subsets of input tokens. Visual tokens from each frame are randomly partitioned into two spatial groups, with the number of tokens in the first group uniformly sampled. Attention masks within transformer layers constrain interactions to occur only among tokens within the same spatial group, and this grouping remains consistent across time steps. Consequently, the model becomes capable of processing varying subsets of visual tokens effectively, while optimizing the same prediction objective. We demonstrate its necessity in the ablation studies.
Pointmaze (60 evals) Wall (60 evals) PushT (60 evals) Granular (20 evals) Rope (30 evals) Block Pushing (50 evals)
Full 98.3 85.0 75.0 75.0 63.3 22.0
CLS 96.7 91.7 43.3 20.0 36.7 16.0
Drop (Ratio)
10% 91.7 93.3 78.3 80.0 73.3 28.0
20% 90.0 95.0 70.0 80.0 66.7 16.0
30% 98.3 93.3 61.7 85.0 70.0 18.0
40% 96.7 93.3 56.7 70.0 76.7 18.0
50% 100.0 95.0 70.0 60.0 73.3 20.0
60% 83.3 91.7 46.7 70.0 70.0 20.0
70% 70.0 93.3 28.3 55.0 53.3 20.0
80% 73.3 90.0 21.7 50.0 66.7 24.0
90% 71.7 86.7 20.0 40.0 50.0 12.0
Pointmaze Wall PushT Block Pushing
Planning Time (s/iter) Change (%) Planning Time (s/iter) Change (%) Planning Time (s/iter) Change (%) Planning Time (s/iter) Change (%)
Full 184 0.0 79 0.0 173 0.0 297 0.0
CLS 49 -73.4 40 -49.4 32 -81.5 163 -45.1
Drop (Ratio)
10% 165 -10.3 73 -7.6 149 -13.9 278 -6.4
20% 141 -23.4 69 -12.7 131 -24.3 259 -12.8
30% 126 -31.5 65 -17.7 114 -34.1 240 -19.2
40% 106 -42.4 62 -21.5 97 -43.9 214 -27.9
50% 93 -49.5 53 -32.9 82 -52.6 208 -30.0
60% 80 -56.5 50 -36.7 69 -60.1 200 -32.7
70% 69 -62.5 46 -41.8 59 -65.9 184 -38.0
80% 56 -69.6 46 -41.8 49 -71.7 175 -41.1
90% 48 -73.9 42 -46.8 38 -78.0 167 -43.8
4 Experiments and Analysis
@@ 4.1 Implementation Details
We conduct our evaluations on six datasets (See Fig. 3), each designed to test different aspects of visual planning: maze-based navigation (Pointmaze, Wall), tabletop object pushing (PushT), manipulation of deformable objects using robotic arms (Rope, Granular), and precise robotic manipulation tasks (Block Pushing). In each environment, the goal is reaching a target observation randomly sampled as a desired goal state, starting from initial positions. All observations consist of RGB images with a resolution of $224\times 224$.
Figure 3: Environments. We assess the world model across six distinct environments, arranged as follows (left-to-right, top-to-bottom): Pointmaze, Wall, PushT, Granular, Rope, and Block Pushing.
@@
We evaluate our approach across several visual control domains and compare it against key baselines. The primary baselines include: (1) a Full-Patch world model, which shares the same architecture as ours but utilizes all patches during planning (i.e., $p=0$) and applies positional encoding to each patch—representing a high-fidelity yet computationally expensive setting; and (2) a CLS-Token world model, which employs only the global CLS token from the DINO encoder as the visual representation, thereby offering a single-vector representation baseline. These two baselines are based on the approach presented in zhou2024dinowm . Our analysis focuses on the trade-offs between these approaches in terms of task performance and inference efficiency.
@@
We apply MPC-CEM to the PointMaze, Wall, PushT, and Block Pushing environments, whereas the Granular and Rope environments employ the CEM approach due to the prolonged processing time in the deformable environment. Here, CEM refers to optimizing a fixed sequence of actions without intermediate corrections or replanning during execution, whereas MPC-CEM denotes optimization using CEM with replanning at each step in a receding horizon manner. In all environments, the MPC horizon is set to $H=5$. For each optimization step, we sample 100 candidate actions and select the top 10 based on the distance computed between the current and goal observations. Note that in sparse imagination scenarios, the distance is computed using only a subset of patches rather than the full patches. Each CEM optimization is performed for 10 steps. For the Wall and Block Pushing environments, the maximum number of MPC iterations is set to 10, while for Pointmaze and PushT, we set the maximum iterations to 15.
4.2 Performance and Computational Efficiency
The experimental outcomes are summarized across two primary tables. Table 1 reports the Success Rate (SR) to the goal across various environments with various token dropout ratios. We evaluated 60 episodes for PointMaze, Wall, and PushT, and 50 episodes for Block Pushing. Due to the prolonged interaction time in the Deformable Environment, Rope and Granular are evaluated on 30 and 20 episodes, respectively. Table 2 outlines the associated planning times, highlighting percentage reductions compared to the Full-Patch model. Employing moderate dropout ratios (typically between 10% and 50%) yields comparable or improved performance relative to the Full-Patch baseline, with substantial reductions in planning times. For example, around a dropout level 50%, planning time reductions approach 30–50% in tested environments. Although dropout models remain slower than the CLS-Token baseline, their predictive performance significantly surpasses the CLS-based method.
Task Difficulty and Scene Complexity.
This tendency is more pronounced in complex or difficult environments (Granular, Rope, PushT, Block Pushing). In spatially demanding scenes (Granular, Rope), the CLS-Token baseline (20.0%, 36.7% success) fails to capture necessary details. In contrast, our patch dropout approach with moderate ratios achieves significantly higher success (up to 85.0%, 76.7%), demonstrating that a pruned token set better preserves critical spatial information than a single global token. For high-difficulty tasks (PushT, Block Pushing), dropout also substantially outperforms CLS (43.3%, 16.0%) and performs comparably to or slightly better than the Full-Token baseline (e.g., 75.0% for Full in PushT), validating the efficacy of planning with token subsets.
Dropout Ratios.
Generally, performance remains stable compared to using full patches for planning, with dropout ratios up to approximately 50%, after which it deteriorates, although the degree of degradation varies by dataset. This suggests that excessively high dropout rates result in the loss of patches containing critical information required for effective task completion.
Planning Time Improvements.
In addition to improved performance, our patch dropout method offers significant computational benefits. For instance, in the PushT environment, planning time per iteration drops from 173 seconds (Full-Patch baseline) to 82 seconds at a 50% dropout ratio—52.6% reduction. Similar improvements in planning time are observed across other environments such as Wall, PushT, and Block Pushing. Although the CLS-Token baseline provides even faster inference times, its dramatic drop in performance—especially in complex and high-difficulty scenarios—renders it less viable for practical applications.
Figure 5: nHSIC between visual tokens and environment states. nHSIC scores with respect to dropout rate for random dropout compared to Full, CLS, and Attention-Encoder dropout in Granular dataset.
Figure 6: Attentive probing performance. Validation losses for state prediction with attentive probing compared to Attention-Encoder baseline in Granular dataset.
4.3 Empirical Insights into Sparse Imagination
We validate our approach by showing experimentally that a subset of visual patch tokens is sufficient for effective planning, making the full patch tokens unnecessary.
Information-Theoretic Analysis.
We measured the mutual information between visual tokens and ground-truth states using normalized Hilbert-Schmidt Independence Criterion (nHSIC) 3437 to assess token dropout’s impact on information quality. Experiments across various dropout rates (Fig. 5) show that high HSIC scores are maintained even under substantial dropout, indicating minimal loss of critical state information. Furthermore, even retaining just 10% of tokens yields higher nHSIC than the CLS token, which is intended to capture global image information.
Attentive Probing Analysis.
To further validate our findings, we trained an attentive probing module bardes2024revisiting to assess how well tokens predict ground-truth states (details in Fig. 6 and Appendix). While probing errors modestly increase with higher dropout rates, this suggests random subsets still encode sufficient information. Furthermore, a single random token shows predictive accuracy comparable (albeit with variance) to the CLS token, indicating state information is distributed across tokens.
Role of Randomized Grouped Attention.
While a subset of patch tokens contains sufficient state information for accelerated planning via token dropout, predicting solely on this subset can introduce errors. We show that training the world model with randomized grouped attention can reduce these errors. Our comparison of full attention and grouped attention models shows that grouped attention lowers the mean relative L2 distance of subset predictions (0.016) compared to full attention (0.036). This reduction in prediction error aligns with the improved planning performance observed with grouped attention under token dropout (Fig. 4). Experimental details on attention mechanisms are provided in the Appendix.
Planning Robustness.
Although randomized grouped attention reduces prediction errors on the token subset under dropout, complete elimination is not achieved. Consistent with prior studies showing CEM’s robustness to noise and uncertainties de2005tutorial ; hu2009performance ; lale2024falcon , we claim this remaining error sufficient for effective planning, based on the experimental results (Table 1). Detailed experiments evaluating CEM’s performance with visual token errors in our framework are in the Appendix.
4.4 Token Dropout Methods
We investigated whether more sophisticated token dropout strategies can improve planning performance. To this end, we empirically evaluated the average planning success rate of various token selection and merging methods across different drop ratios. We compare uniform baselines against more advanced learned or structured approaches. As uniform baselines within the MPC framework, we consider Random (our standard setting), which resamples the dropout mask at each MPC step, and Fixed, which uses a single mask throughout. Both sample tokens uniformly. Beyond these, we evaluate:
  • LHS mckay2000comparison : Employs Latin Hypercube Sampling for spatially uniform sampling at each step.
  • LTRP luo2024ltrp : Trains a score network to predict patch importance using the reconstruction loss from pre-trained MAE he2022mae .
  • Attention-Encoder: Derives token importance from the attention map between the CLS and patch tokens of a DINO encoder, where higher attention weights indicate greater importance. This is a common practice in pruning-based prior research Liang2021evit ; wang2022vtclfc ; Zhang2023star . We use attention map of the final layer of pretrained DINO.
  • STAR Zhang2023star : Calculates importance scores by fusing intra-layer scores from CLS attention and inter-layer scores from modified Layer-wise Relevance Propagation (LRP; chefer2021transformer ). We select important tokens based on the final layer score of DINO encoder.
  • Attention-WM: Evaluates importance based on the average attention weight a token receives as a query from others in each layer. Importance is determined by averaging the attention maps across all layers.
  • Agglomerative Token Clustering (ATC) haurum2024agglomerative : Utilizes agglomerative clustering to group tokens, followed by average pooling. For planning, we first apply agglomerative clustering to the context observation image features to predict future clustered features. In parallel, agglomerative clustering is performed on the goal image features. The distance between the predicted future and the goal is then determined via Hungarian matching between the predicted future clusters and the goal clusters, and this distance guides the planning process.
Table 3: The average planning success rate (%) achieved by various token dropout and merging methods under different drop ratios. Our results suggest that no method achieves significantly superior performance compared to the Random sampling.

### v2

@@ Abstract
World model based planning has significantly improved decision-making in complex environments by enabling agents to simulate future states and make informed choices. This computational burden is particularly restrictive in robotics, where resources are severely constrained. To address this limitation, we propose a Sparse Imagination for Efficient Visual World Model Planning, which enhances computational efficiency by reducing the number of tokens processed during forward prediction. Our method leverages a sparsely trained vision-based world model based on transformers with randomized grouped attention strategy, allowing the model to flexibly adjust the number of tokens processed based on the computational resource. By enabling sparse imagination during latent rollout, our approach significantly accelerates planning while maintaining high control fidelity. Experimental results demonstrate that sparse imagination preserves task performance while dramatically improving inference efficiency. This general technique for visual planning is applicable from simple test-time trajectory optimization to complex real-world tasks with the latest VLAs, enabling the deployment of world models in real-time scenarios.
@@ 1 Introduction
By “imagining” future trajectories in a learned world model of the environment, agents can perform sophisticated decision making without trial-and-error in the real world (Ha and Schmidhuber, 2018; Hafner et al., 2019a; b). Recent world model advancements for visual control tasks, which necessitate reasoning over high-dimensional image observations, increasingly leverage high-level visual features derived from self-supervised vision models (Caron et al., 2021; Oquab et al., 2023; Nair et al., 2022; Xiao et al., 2022). For example, DINO-WM (Zhou et al., 2024) introduces a world model that predicts future Vision Transformer (ViT) patch tokens (DINO features) instead of pixels or single vector representation, enabling zero-shot planning for diverse tasks. These approaches retain rich spatial tokens, achieving strong generalization on complex control tasks. However, using numerous visual tokens for planning incurs a prohibitive quadratic computational cost, as it requires many simulated rollouts. This raises a crucial question:
@@
In this paper, we address the above question by introducing the concept of sparse imagination, accelerating world model inference by deliberately using only a sparse subset of visual tokens during forward prediction. Our key insight is that ViT-based representations are redundant (Raghu et al., 2021; Pan et al., 2021; Chen et al., 2022; Kim et al., 2024; Siméoni et al., 2025); thus, not all patch tokens are necessary for planning.
@@
  • We introduce sparse imagination, a simple yet effective method that enables efficient visual world model planning by leveraging random patch feature dropout during inference.
  • We demonstrate that our approach is a general technique applicable from simple test-time trajectory optimization to complex real-world tasks with VLAs. It achieves substantial planning speedups while maintaining high performance, validated on benchmarks like LIBERO-10 (Liu et al., 2023a), Meta-World (Yu et al., 2021), and on real-world robotic tasks.
  • Our comparative analysis shows that simple random sampling achieves competitive or superior performance to sophisticated token selection methods due to a fundamental “blind spot” problem we identify, where static importance metrics fail in dynamic planning, making the unbiased coverage of random sampling a more robust strategy.
@@ 2 Related Work
Early world models transitioned from computationally heavy pixel-level models (Finn et al., 2016; Ebert et al., 2018) to latent dynamics models like PlaNet (Hafner et al., 2019b) and Dreamer (Hafner et al., 2019a; 2020; 2023), which achieved impressive results by leveraging planning and latent imagination. While most existing latent world models compress entire images into relatively low-dimensional vectors, these compact representations often lose the fine-grained spatial information necessary for high-precision manipulation tasks (Zhou et al., 2024; Tsagkas et al., 2025).
@@
A recent trend in decision making leverages powerful pre-trained visual encoders for rich state representations (Nair et al., 2022; Xiao et al., 2022). While effective for representation learning, methods using a single image embedding (e.g., CLS token) from encoders like ResNet or ViT often lose crucial fine-grained spatial details, hindering performance on tasks requiring precise spatial understanding, such as manipulation (Tsagkas et al., 2025; Zhou et al., 2024). To address this, approaches using spatial features like ViT patch embeddings for decision making have been developed, exemplified by DINO-WM (Zhou et al., 2024) which uses a set of DINO patch tokens as the state. Building on this paradigm, our work utilizes pre-trained ViT patch features for their spatial granularity. Our innovation is an inference-time token dropping mechanism that significantly enhances computational efficiency during imagination.
@@
The quadratic scaling of ViT self-attention poses a significant computational challenge, particularly for sequential models like world models (Dosovitskiy et al., 2020). Improving inference efficiency via token reduction is thus a key vision research area. Existing methods employ diverse strategies such as learned/attention-based selection (Rao et al., 2021; Luo et al., 2024; Yin et al., 2022; Liang et al., 2021; Wang et al., 2022; Zhang et al., 2024b), merging (Bolya et al., 2022; Feng and Zhang, 2023; Haurum et al., 2024), and training-time dropout (Liu et al., 2023b). These demonstrate efficiency gains from ViT representation redundancy (Raghu et al., 2021; Pan et al., 2021; Chen et al., 2022; Kim et al., 2024). In contrast, we present a patch-dropping method specifically for inference-time world model planning. Our approach uniquely applies dropout only during the imagination phase, enabling seamless integration with any transformer world model.
@@ 3 Methodology
Figure 1: Sparse Imagination. Sparse imagination accelerates planning by performing model predictive control (MPC) rollouts on a random subset of visual tokens. A new dropout pattern is dynamically sampled at each MPC iteration, and both predictions and optimization for CEM are computed using only the selected patches to improve efficiency and robustness.
@@ 3.2 Sparse Imagination for MPC
We use MPC for planning, and for each step, candidate action sequences $\{a_{t:t+H-1}\}$ are sampled from a distribution (CEM) or pre-trained policy. The trained world model $f_{\theta}$ then predicts future states by iteratively rolling out these action sequences from the current state. In a goal-conditioned setting, we evaluate candidate sequences using the MSE loss between the predicted and goal state features:
[…]
@@
When using CEM for test-time trajectory optimization, the distribution of candidate actions is subsequently updated based on the top-performing sequences. This iterative optimization continues for a predefined iteration step $M$, after which the optimized action sequence is executed in the environment.
@@
MPC is computationally intensive, as world model rollout occurs $K\times M\times H$ per every planning step, and this cost scales quadratically with the number of visual tokens (the number of candidate sequences is $K$ and planning horizon is $H$), making real-time execution increasingly challenging. Since attention layers in the world model are the primary source of this computational burden, reducing the number of tokens processed is crucial for improving planning efficiency.
@@
We address this problem through sparse imagination. Given that our model is explicitly trained to operate on arbitrary token subsets, we implement random token dropout during planning as shown in Fig. 1. At each planning step, we randomly generate a token dropout mask characterized by a dropout fraction $p\in[0,1)$ of the user’s choice, denoting the proportion of tokens to discard, keeping randomly sampled (1p)N(1-p)N tokens. Sparse rollout imagination can effectively reduce redundant computations and improve planning efficiency. Planning a single step might fall to wrong direction when dropping key features but has minimal impact on task performance since it can recover from failures by re-sampling at each planning iteration. The drop ratio $p$ serves as a controllable parameter balancing computational speed and task accuracy. We empirically evaluate and analyze this trade-off in the following Sections.
Figure 2: Randomized Grouped Attention Strategy. Our randomized grouping strategy used during training to generalize to arbitrary token subsets. Visual tokens are randomly partitioned into two groups, and attention is masked to occur only within the same spatial group. This trains the model to process sparse inputs effectively, and its necessity is shown in ablations.
4 Experiments
@@ 4.1 Implementation Details
We evaluate our method on eight simulated environment and two real-world tasks using LeRobot (Cadene et al., 2024) (See Fig. 3). Detailed descriptions of the environments, datasets, and model configurations are provided in the Appendix A.
Figure 3: Environments. We assess the world model across eight simulation environments and two real-world tasks, arranged as follows (left-to-right, top-to-bottom): Pointmaze, Wall, PushT, Granular, Rope, Block Pushing, LIBERO, Meta-World and LeRobot tasks (PickPlace & Drawer).
@@
In each environment, the goal is reaching a target observation randomly sampled as a desired goal state, starting from initial positions. For the LIBERO environment, we conducted experiments on LIBERO-10, a suite of long-horizon tasks, to demonstrate the effectiveness of our planning acceleration. For the real world, the task is to pick the blue block and place it in the cup.
@@
We evaluate our approach across several visual control domains and compare it against key baselines. The primary baselines include: (1) a Full-Patch world model, DINO-WM (Zhou et al., 2024) based on DINO (Caron et al., 2021), which shares the same architecture as ours but utilizes all patches during planning (i.e., $p=0$) and applies positional encoding to each patch—representing a high-fidelity yet computationally expensive setting; and (2) a CLS-Token world model, which employs only the global CLS token from the DINO encoder as the visual representation, thereby offering a single-vector representation baseline. These two baselines are based on the approach presented in DINO-WM.
@@
To evaluate our framework across diverse scenarios, we utilize three distinct planning schemes:
  • MPC-CEM: For Pointmaze, Wall, PushT, Block Pushing, we use MPC-CEM. This involves replanning at each step in a receding-horizon manner.
  • CEM (Open-loop): For computationally intensive deformable object simulations (Granular, Rope), we employ a standard open-loop CEM approach. This optimizes a single, fixed sequence of actions without replanning to manage the high processing cost.
  • Policy-Guided Planning: For complex, long-horizon tasks (LIBERO, Meta-World, LeRobot), where random action sampling is inefficient, we guide the planner by sampling $K$ candidate action sequences from a pre-trained policy such as Vision-Language-Action (VLA) models. Our planner then evaluates these candidates.
Our contribution highlights significant efficiency gains while preserving Full-Patch (DINO-WM) baseline performance, demonstrating our sparse imagination as an effective, computationally efficient replacement, rather than focusing on absolute performance maximization. Therefore, for fair comparison, most of hyperparameters are adopted from the DINO-WM baseline, as detailed in the Appendix A.3 and B.1.